OpenStudy (anonymous):

what is the distance from AB when AC is 6 feet and CB is 8 feet?

OpenStudy (anonymous):



OpenStudy (anonymous):

Have you heard of the Pythagorean Theorem?

OpenStudy (anonymous):

 Pythagorean Theorem:\[\large a^2 + b^2 = c^2\]C always stands for the hypotenuse.

OpenStudy (anonymous):

What is the length of side A B? ten feet eleven feet twelve feet thirteen feet

OpenStudy (anonymous):

\[\large a^2 + b^2 = c^2\]\[\large 6^2 + 8^2 = c^2\]\[\large 36 + 64 = c^2\]\[\large 100 = c^2\]Find the square root to solve for c.\[\sqrt{100} =\sqrt{c^2}\]\[\large 10 = c\]

OpenStudy (anonymous):

@yoyogators

OpenStudy (anonymous):

okay I understand